if you add me to myself and then add 20, you get 36, which number i am.
Found 2 solutions by Pushkar, MathTherapy:
Answer by Pushkar(3)

See:-
36 - 20 = 16
So 8 + 8 = 116
The number is 8
As the question:- if you add me to myself. The both numbers will be same. So, your answer is 8.

Answer by MathTherapy(10552)


if you add me to myself and then add 20, you get 36, which number i am.

Let number be N
Then, N + N + 20 = 36
Solve for N
OR
Work backwards
1) Subtract 20 from 36
2) Divide result in 1) by 2